### v2

# Heegaard Floer Surgery Formula and Cosmetic Surgeries

Abstract
Two Dehn surgeries on a knot are called cosmetic if they yield homeomorphic three-manifolds. We show for a certain family of null-homologous knots in any closed orientable three-manifold, if the knot admits cosmetic surgeries with a pair of positive surgery coefficients, then the coefficients are both greater than $1$. In addition, for this family of knots, we show that +1/q+1/q Dehn surgery is not homeomorphic to the original three-manifold. The proofs of these results use the mapping cone formula for the Heegaard Floer homology of Dehn surgery in terms of the knot Floer homology of the knot; we provide a new proof of this formula for integer surgeries in $\operatorname{Spin}^{c}$ structures with nontorsion first Chern class.
Keywords: Heegaard Floer Homology, knot Floer homology, Dehn surgery, cosmetic surgery
1 Introduction
Heegaard Floer homology is a powerful invariant of closed oriented three-manifolds introduced by Ozsváth and Szabó [OS04b]. It has been used to prove many results about three-manifold topology, especially Dehn surgery. In fact, the Heegaard Floer homology of any surgery on a knot can be computed from the knot Floer complex of the knot [OS08, OS10, ZEM23].
In this paper, we will use Heegaard Floer homology to study cosmetic surgeries on a certain class of knots in arbitrary closed oriented three-manifolds. We recall the definition of cosmetic surgeries.
Definition 1.1.
If two Dehn surgeries on a knot yield homeomorphic manifolds, then these two surgeries are cosmetic. If in addition, the two surgeries are orientation-preserving homeomorphic, then these two surgeries are purely cosmetic.
Cosmetic surgeries are very rare; in fact, the following cosmetic surgery conjecture has been proposed:
Conjecture 1.2 ([KIR95] Problem 1.81).
Suppose $K$ is a knot in a closed oriented three-manifold $Y$ such that YKY-K is irreducible and not homeomorphic to the solid torus. If two different Dehn surgeries on $K$ are purely cosmetic, then there is a homeomorphism of YKY-K which takes one slope to the other.
The existence of cosmetic surgeries has been ruled out in many cases for knots in $S^{3}$. For instance, Ozsváth and Szabó proved that if $K\subset S^{3}$ admits cosmetic rational surgeries, then either the resulting manifold is an $L$-space or the surgery slopes have opposite signs [OS10]. Much more is known when considering purely cosmetic surgeries. Hanselman proved that any pair of purely cosmetic Dehn surgeries on knots in $S^{3}$ must have slopes that are $\pm 2$ or $\pm 1/q$ for some integer $q$ [HAN19].
Less is known for knots in arbitrary three-manifolds. Ni found an obstruction to cosmetic surgeries on knots in three-manifolds satisfying a certain condition on their Thurston norms [NI10].
The main theorem in this paper is an analogue of Ozsváth and Szabó’s result. We consider a class of knots which were shown by Ni to have special properties on their knot Floer homology.
Definition 1.3.
Let $Y$ be a closed, oriented three-manifold, and $K\subset Y$ a null-homologous knot. We say that $K$ is blow-up bounding if there exists a taut Seifert surface $F$ for $K$, some natural number $N$, and a properly-embedded smooth disk $G$ inside $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ with $\partial G=K\times\{0\}$ and $G]=\iota_{*}[F]\in H_{2}(X,K\times\{0\})$ for $\iota:(Y,K)\to(X,K\times\{0\})$ the inclusion.
We say that $K$ is blow-up unknotted if $K$ is obtained from the unknot in $Y$ by a sequence of crossing changes that change a positive crossing to a negative one.
By Kirby calculus, $K$ is blow-up bounding if $K$ is blow-up unknotted.
Remark.
The property of $K$ being blow-up bounding implies that $K$ is null homotopic.
Theorem 1.4.
Let $K\subset Y$ be a blow-up unknotted knot, and suppose $0<r<s$ are distinct positive rational numbers with the property that $Y_{r}(K)\cong\pm Y_{s}(K)$. Then $r,s>1$ or $K$ is the unknot.
We also provide progress on the knot complement problem for blow-up unknotted knots.
Theorem 1.5.
Let $K\subset Y$ be a nontrivial blow-up bounding knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $K$ is a nontrivial blow-up unknotted knot, then $Y_{1}(K)\ncong\pm Y$.
The proofs of these theorems rely on a mapping cone formula that relates the Heegaard Floer homology of Dehn surgery on null-homologous knots in arbitrary closed, oriented three-manifolds with the knot Floer chain complex. This result was proven for Heegaard Floer homology in $\operatorname{Spin}^{c}$ structures whose first Chern classes are torsion by Ozsváth and Szabó, and the formula in nontorsion $\operatorname{Spin}^{c}$ structures for zero surgeries was proven by Ni [OS08, OS10, NI20]. A recent result by Zemke provides a very general surgery formula via an equivalence of certain Lagrangians with local systems in the Fukaya category [ZEM23]. Perhaps of independent interest, we provide a conceptually-simpler proof of the integer surgery formula for nontorsion $\operatorname{Spin}^{c}$ structures, which generalizes Ozsváth and Szabó’s and Ni’s results.
2 Preliminaries on Heegaard Floer Homology
In this section, we establish some notation and collect some results, most of which are copied from [OS08], that will be needed in later sections.
Let $W:Y_{1}\to Y_{2}$ be an oriented cobordism, with $Y_{1},Y_{2}$ connected. Given $\mathfrak{s}\in\operatorname{Spin}^{c}(W)$, there is a homomorphism
$F_{W,\mathfrak{s}}^{\circ}:HF^{\circ}(Y_{1},\mathfrak{s}|Y_{1})\to HF^{\circ}(Y_{2},\mathfrak{s}|Y_{2}),$
induced by a chain map
$f_{W,\mathfrak{s}}^{\circ}:CF^{\circ}(Y_{1},\mathfrak{s}|Y_{1})\to CF^{\circ}(Y_{2},\mathfrak{s}|Y_{2}).$
Among other properties, these maps satisfy a conjugation invariance. For $\mathfrak{t}$ a $\operatorname{Spin}^{c}$ structure (in either three or four dimensions), let $\overline{\mathfrak{t}}$ denote its conjugate. There is an isomorphism $\mathfrak{I}_{Y}:HF^{\circ}(Y,\mathfrak{t})\to HF^{\circ}(Y,\overline{\mathfrak{t}})$.
Theorem 2.1 ([OS06] Theorem 3.6).
Let $W$ be a cobordism from $Y_{1}$ to $Y_{2}$ as above. Then
$F_{W,\mathfrak{s}}^{\circ}=\mathfrak{I}_{Y_{2}}\circ F_{W,\overline{\mathfrak{s}}}^{\circ}\circ\mathfrak{I}_{Y_{1}}.$
Given a chain map $f:A\to B$, let $MC(f)$ be the mapping cone of $f$, namely, $MC(f)=A\oplus B$ with the differential
$\partial=\begin{pmatrix}\partial_{A}&0\\ f&\partial_{B}\end{pmatrix}$.
Theorem 2.2 ([OS08] Theorem 3.1).
Let $Y$ be a closed, oriented three-manifold, $K\subset Y$ a null-homologous knot. Fix an integer $n$ and a positive integer $m$. Then, there is a long exact sequence
$\cdots\to HF^{+}(Y_{n}(K))\to HF^{+}(Y_{m+n}(K))\to\bigoplus^{m}HF^{+}(Y)\to\cdots$
and also a corresponding exact sequence using $\widehat{HF}$ in place of $HF^{+}$. There are ${\mathbb{Z}}[U]$-equivariant chain maps
$f_{1}^{+}:CF^{+}(Y_{n}(K))\to CF^{+}(Y_{m+n}(K))$
$f_{2}^{+}:CF^{+}(Y_{m+n}(K))\to\bigoplus^{m}CF^{+}(Y)$
$f_{3}^{+}:\bigoplus^{m}CF^{+}(Y)\to CF^{+}(Y_{n}(K))$
inducing the maps in the long exact sequence, and ${\mathbb{Z}}[U]$-equivariant quasi isomorphisms
$\phi^{+}:CF^{+}(Y_{n}(K))\to MC(f_{2}^{+})$
$\psi^{+}:MC(f_{2}^{+})\to CF^{+}(Y_{n}(K)).$
[…]
Write the two factors of the $\mathbb{Z}$$\oplus{\mathbb{Z}}$ filtration as $(i,j)$. There is a $U$-action on $C$ that is an isomorphism of bifiltration (1,1)(-1,-1).
We will often use the following fact, for example see [HW18, Reduction Lemma].
Lemma 2.3.
With coefficients in any field $\mathbb{F}$}, the knot Floer complex $(C,\partial)$ for any knot $K\subset Y$ is $({\mathbb{Z}}\oplus{\mathbb{Z}})$-filtered chain homotopy equivalent to a chain complex $(\overline{C},\overline{\partial})$ such that the restriction of the differential $\overline{\partial}$ to any $\overline{C}\{i,j\}$ is 0 and $\overline{C}\{0,j\}$ is isomorphic to $\widehat{HFK}(K,j)$.
Knot Floer homology is known to detect the genus of the knot, and this is true for knots in general three-manifolds.
Theorem 2.4.
[NI10, Theorem 2.4] For any field $\mathbb{F}$}, if $K\subset Y$ is a knot of genus $g$, then
$\widehat{HFK}(Y,K,s;{\mathbb{F}})=0$
if $|s|>g$. Furthermore, $\widehat{HFK}(Y,K,g;{\mathbb{F}})\neq 0$.
Define $A_{s}^{+}=C\{\max(i,j-s)\geq 0\}$ and $B^{+}=C\{i\geq 0\}$; these are chain complexes because the differential decreases both filtrations. There are two canonical chain maps $v_{s}^{+}:A_{s}^{+}\to B^{+}$ and $h_{s}^{+}:A_{s}^{+}\to B^{+}$, where $v_{s}^{+}$ is the vertical projection, and $h_{s}^{+}$ first projects $A_{s}^{+}$ to $C\{j\geq s\}$, then maps to $C\{j\geq 0\}$ via $U^{s}$, and finally maps to $B^{+}$ via a chain homotopy equivalence $C\{j\geq 0\}\to C\{i\geq 0\}$ [OS08].
For all $s\in{\mathbb{Z}}$, there is a chain homotopy equivalence $\phi$ between $A_{s}^{+}$ and $A_{-s}^{+}$: first, the $U^{s}$ map is an isomorphism
$A_{s}^{+}=C\{\max(i,j-s)\geq 0\}\to C\{\max(i,j-s)\geq-s\},$
then the change of basepoint map that swaps the roles of $w$ and $z$ in the doubly-pointed Heegaard diagram is a chain homotopy equivalence
$C\{\max(i,j-s)\geq-s\}\to C\{\max(i,j+s)\geq 0\}=A_{-s}^{+}.$
Let $\phi$ be the composition of these two maps. It is clear from the definitions that the following lemma holds:
Lemma 2.5.
For all $s\in{\mathbb{Z}}$, $v_{s}^{+}$ and $h_{s}^{+}$ are chain homotopy equivalent to $h_{-s}^{+}\circ\phi$ and $v_{-s}^{+}\circ\phi$, respectively.
Since all maps involved are $U$-equivariant, the above lemma also holds for the $CF^{\delta}$ version of knot Floer homology for any $\delta\geq 0$.
Lemma 2.6.
Let $g$ be the genus of $K$, then for all $\delta\geq 0$ and $s\geq g+\delta$, $v_{s}^{\delta}$, and $h_{-s}^{\delta}$ induce isomorphisms on homology, and $v_{-s-1}^{\delta}$ and $h_{s+1}^{\delta}$ are nullhomotopic.
Proof.
It is shown in the proof of the adjunction inequality in [OS04a, Theorem 5.1] that one can construct a special Heegaard diagram for $(Y,K)$ such that all generators of $\widehat{CFK}(Y,K)$ lie in Alexander grading greater than or equal to g-g. Thus, all generators for $A_{s}^{\delta}$ with $i\geq 0$ must have $j\geq-g$, from which it follows that $h_{-s}^{\delta}$ induces an isomorphism on homology and $v_{-s-1}^{\delta}$ is null-homotopic. By the $CF^{\delta}$ version of Lemma 2.5, we obtain the statements about $v_{s}^{\delta}$ and $h_{s+1}^{\delta}$. ∎
Let $K\subset Y$ be null-homologous. Let $F$ be a Seifert surface for $K$ and let $\hat{F}$ be the closed surface obtained by capping off $\partial F$ with a disk. Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, which need not have torsion first Chern class, and let $n\in{\mathbb{Z}}$. The space of $\operatorname{Spin}^{c}$ structures over $Y_{n}(K)$ which are $\operatorname{Spin}^{c}$-cobordant to $\mathfrak{t}$ over the natural two-handle cobordism $W_{n}^{\prime}(K):Y_{n}(K)\to Y$ is identified with ${\mathbb{Z}}/n{\mathbb{Z}}$ by $\mathfrak{u}\mapsto i$, where $\mathfrak{s}$ is an extension of $\mathfrak{u}$ over $W_{n}^{\prime}(K)$ with the property that
$\langle c_{1}(\mathfrak{s}),[\hat{F}]\rangle-n\equiv 2i\pmod{2n}$.
For $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ denote the corresponding summand of $CF^{+}(Y_{n}(K))$. Let $A_{s,\mathfrak{t}}^{+},B_{\mathfrak{t}}^{+}$ denote the corresponding chain complexes associated to the knot filtration on $CF^{+}(Y,\mathfrak{t})$ [OS08].
Let
${\mathbb{A}}_{i,\mathfrak{t}}^{+}=\bigoplus_{s\in{\mathbb{Z}}|s\equiv i\pmod{n}}A_{s,\mathfrak{t}}^{+}\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{\mathfrak{t}}^{+}=\bigoplus_{s\in{\mathbb{Z}}|s\equiv i\pmod{n}}B_{\mathfrak{t}}^{+}$
and let $D_{n,i,\mathfrak{t}}^{+}:{\mathbb{A}}_{i,\mathfrak{t}}^{+}\to{\mathbb{B}}_{i,\mathfrak{t}}^{+}$ be the map
$D_{n,i,\mathfrak{t}}^{+}(\{a_{s}\})=\{b_{s}\},$
where here
$b_{s}=h_{s-n}^{+}(a_{s-n})+v_{s}^{+}(a_{s}).$
Let ${\mathbb{X}}_{i,\mathfrak{t}}^{+}(n)$ denote the mapping cone of $D_{n,i,\mathfrak{t}}^{+}$; we prove in the next section that this chain complex computes $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ [OS08].
Taking the direct sum of the above chain complexes over all $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ and $i\in{\mathbb{Z}}/n{\mathbb{Z}}$ gives $\mathbb{A}$$^{+}$, $\mathbb{B}$$^{+}$, and the map $D_{n}^{+}:{\mathbb{A}}^{+}\to{\mathbb{B}}^{+}$. Let ${\mathbb{X}}^{+}(n)$ denote the mapping cone of $D_{n}^{+}$.
Let $K\subset Y$ be a null-homologous knot and let $F$ be a Seifert surface for $K$ in the homology class $\phi$. Given $n>0$ and $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $\widehat{F}_{n}\subset W_{n}^{\prime}(K)$ be the closed surface obtained by capping off $\partial F$ with a disk. Let $\widehat{\phi}_{n}=[\widehat{F}_{n}]\in H_{2}(W_{n}^{\prime}(K))$. Let $k$ be an integer satisfying $k\equiv i\pmod{n}$ and $|k|\leq n/2$. Let $\mathfrak{x}_{k},\mathfrak{y}_{k}\in\operatorname{Spin}^{c}(W_{n}^{\prime}(K))$ satisfy $\mathfrak{x}_{k}|Y=\mathfrak{y}_{k}|Y=\mathfrak{t}$ and
$\langle c_{1}(\mathfrak{x}_{k}),\hat{\phi}_{n}\rangle=2k-n,\hskip 14.22636pt\langle c_{1}(\mathfrak{y}_{k}),\hat{\phi}_{n}\rangle=2k+n,$
and let $\mathfrak{t}_{i}=\mathfrak{x}_{k}|Y_{n}(K)=\mathfrak{y}_{k}|Y_{n}(K)$.
Theorem 2.7 ([OS08] Theorem 2.3).
When $n\geq 2g(F)$, the chain complex $CF^{+}(Y_{n}(K),\mathfrak{t}_{i})$ is represented by the chain complex $A_{k,\mathfrak{t}}^{+}$, in the sense that there is an isomorphism
$\Psi_{n}^{+}:CF^{+}(Y_{n}(K),\mathfrak{t}_{i})\to A_{k,\mathfrak{t}}^{+}.$
Moreover, the squares in the diagrams of Figure 1 commute.
[…]
Figure 1: Commutative diagrams of the maps $v_{k}^{+}$ and $h_{k}^{+}$
The $CF^{\delta}$ version of the theorem holds with the same proof.
In the case of rational surgeries, let $p,q$ be a pair of relatively prime integers. For $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ and $i\in{\mathbb{Z}}/p{\mathbb{Z}}$, denote
${\mathbb{A}}_{p/q,i,\mathfrak{t}}^{+}=\bigoplus_{r\in{\mathbb{Z}}}(r,A_{\lfloor(i+pr)/q\rfloor,\mathfrak{t}}^{+}(K))\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{p/q,i,\mathfrak{t}}^{+}=\bigoplus_{r\in{\mathbb{Z}}}(r,B_{\mathfrak{t}}^{+}).$
Let $D_{p/q,i,\mathfrak{t}}^{+}:{\mathbb{A}}_{p/q,i,\mathfrak{t}}^{+}\to{\mathbb{B}}_{p/q,i,\mathfrak{t}}^{+}$ be the map
$D_{p/q,i,\mathfrak{t}}^{+}\{(r,a_{r})\}=\{(r,b_{r})\},$
where
$b_{r}=v_{\lfloor(i+pr)/q\rfloor}^{+}(a_{r})+h_{\lfloor(i+p(r-1))/q\rfloor}^{+}(a_{r-1}).$
Let ${\mathbb{X}}_{p/q,i,\mathfrak{t}}^{+}$ denote the mapping cone of $D_{p/q,i,\mathfrak{t}}^{+}$; in section 4, we prove that the homology of this chain complex is isomorphic to $HF^{+}(Y_{p/q}(K),i,\mathfrak{t})$ [OS10].
As before, direct summing over all $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ and $i\in{\mathbb{Z}}/p{\mathbb{Z}}$ gives ${\mathbb{A}}_{p/q}^{+}$, ${\mathbb{B}}_{p/q}^{+}$, and $D_{p/q}^{+}:{\mathbb{A}}_{p/q}^{+}\to{\mathbb{B}}_{p/q}^{+}$. Let ${\mathbb{X}}_{p/q}^{+}$ denote the mapping cone of $D_{p/q}^{+}$. In the case where $p/q$ is clear from the context, we will drop the subscript $p/q$ from the notation.
3 Integer Surgeries
In this section, we provide a proof of the mapping cone formula for integer surgeries in the case where the $\operatorname{Spin}^{c}$ structure has nontorsion first Chern class. The formula for torsion $\operatorname{Spin}^{c}$ structures was originally proven by Ozsváth and Szabó [OS10]. Our proof combines ingredients from their argument as well as Ni’s proof of the zero surgery formula.
Let $K\subset Y$ be a null-homologous knot. Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$ with nontorsion first Chern class.
Fix $n>0$. The long exact sequence in Theorem 2.2 decomposes over $\operatorname{Spin}^{c}$ structures on $Y$. Setting $m=nk$ for large $k>0$, it also decomposes over $\operatorname{Spin}^{c}$ structures on $Y_{n}(K)$, i.e. given $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{+})$ where
$f_{2,i,\mathfrak{t}}^{+}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}:CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})\to CF^{+}(Y,\mathfrak{t}),$ (3.1)
$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k+1)}(K))\$$|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k+1)\equiv 2s\pmod{2n(k+1)}$,
$\text{and }s\equiv i\pmod{n}\}$,
$CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ |\ s\equiv i\pmod{n}}CF^{+}(Y_{n(k+1)}(K),s,\mathfrak{t}).$
A strengthening of Proposition 3.5 in [NI20] shows the following:
Proposition 3.2 ([NI20] Proposition 3.5).
Let $n>0,$ $n/2\leq k<n/2$, and $p=\max\{k+g(F),-2k+g(F)\}$. Suppose that $n\geq\max\{p,2g(F)\}$, $\mathfrak{s}\in\mathcal{X}_{k}\setminus\{\mathfrak{x}_{k},\mathfrak{y}_{k}\},$ then the map $F_{W_{n}^{\prime}(K),\mathfrak{s}}^{-}$ factorizes through the map
$U^{n-p}:\operatorname{HF}^{-}(Y_{n}(K),k,\mathfrak{t})\circlearrowright.$
Moreover, if $c_{1}(\mathfrak{t})$ is nontorsion, then $F_{W_{n}^{\prime}(K),\mathfrak{s}}^{+}$ factorizes through
$U^{n-p}:\operatorname{HF}^{+}(Y_{n}(K),k,\mathfrak{t})\circlearrowright.$
This proposition is exactly the same as [NI20] Proposition 3.5 when $k\leq 0$, but when $k>0$, we improve the bound on the power of $U$ from n2kg(F)n-2k-g(F) to np=nkg(F)n-p=n-k-g(F). The proof of the improved bound is just checking that when $k>0$, the inequality
$\frac{l(l+1)n}{2}-lk-g\geq n-2|k|-g$
in Ni’s proof can be improved to
$\frac{l(l+1)n}{2}-lk-g\geq n-k-g.$
By the same proof, the proposition also holds for the $\delta$ version of Heegaard Floer homology for any $\delta\geq 0$.
Theorem 3.3.
For each $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, the mapping cone ${\mathbb{X}}_{i,\mathfrak{t}}^{+}(n)$ of
$D_{n,i,\mathfrak{t}}^{+}:{\mathbb{A}}_{i,\mathfrak{t}}^{+}\to{\mathbb{B}}_{i,\mathfrak{t}}^{+}$
is quasi-isomorphic to $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ as ${\mathbb{Z}}[U]$-modules.
Proof when $n>0$.
We first prove the theorem for the $CF^{\delta}$ version of knot Floer homology for any $\delta\geq 0$, that is, the mapping cone ${\mathbb{X}}_{i,\mathfrak{t}}^{\delta}(n)$ is quasi-isomorphic to $CF^{\delta}(Y_{n}(K),i,\mathfrak{t})$ as ${\mathbb{Z}}[U]$-modules.
Recall that $CF^{\delta}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{\delta})$ as in Equation 3.1. Since $c_{1}(\mathfrak{t})$ is nontorsion, $H_{*}(A_{i}^{+})$ is a finitely generated abelian group. So $U^{q}|H_{*}(A_{i}^{+})=0$ when $q$ is greater than a constant $C_{1}$ independent of $n,k$. Hence, $U^{q}|H_{*}(A_{i}^{\delta})=0$ for $q>C_{1}$ for the same constant $C_{1}$ independent of $n,k,$ and $\delta$. By the $CF^{\delta}$ version of Theorem 2.7, this implies that $U^{q}|HF^{\delta}(Y_{n(k+1)}(K),i,\mathfrak{t})=0$ when $q>C_{1}$ and $n(k+1)>2g$. Choose $k$ such that $n(k+1)>2(g+\delta)$.
Consider a $\operatorname{Spin}^{c}$ structure $\mathfrak{s}_{s}\in\mathcal{X}_{i}$ for some $s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}$. Recall that we choose representatives
$\frac{n(k+1)}{2}\leq s<\frac{n(k+1)}{2}$
for classes in ${\mathbb{Z}}/n(k+1){\mathbb{Z}}$. Suppose $s\geq 0$. Then for $l=\max\{s+g,-2s+g\}$, we have
$l\leq\frac{n(k+1)}{2}+g\leq n(k+1).$
Proposition 3.2 implies that $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}=0$ when $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$.
Now for $s<0$, let $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then by conjugation invariance, there is a chain homotopy equivalence
$F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}\cong F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{\delta}.$
If $\mathfrak{s}_{s}=\mathfrak{x}_{s}+r\operatorname{PD}[\widehat{F}$], then
$\overline{\mathfrak{s}_{s}}=\overline{\mathfrak{x}}_{-s}-(r+1)\operatorname{PD}[\widehat{F}$],
in particular, $\overline{\mathfrak{s}_{s}}$ is not one of the two special $\operatorname{Spin}^{c}$ structures and s>0-s>0. By the above argument, $F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{\delta}=0$, hence also $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}=0$.
Therefore, $MC(f_{2,i,\mathfrak{t}}^{\delta})$ is quasi-isomorphic to the mapping cone of
$\sum_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}\left(f_{W_{n(k+1)}^{\prime}(K),\mathfrak{x}_{s}}^{\delta}+f_{W_{n(k+1)}^{\prime}(K),\mathfrak{y}_{s}}^{\delta}\right):CF^{\delta}(Y_{n(k+1)}(K),[i],\mathfrak{t})\to CF^{\delta}(Y,\mathfrak{t}).$
By Theorem 2.7, this chain complex is quasi-isomorphic to the mapping cone of
$\sum_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}(v_{s}^{\delta}+h_{s}^{\delta}):\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}A_{s}^{\delta}\to\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}B^{\delta}.$
We rewrite this as the truncated mapping cone
$\sum_{-n(k+1)/2\leq s<n(k+1)/2\ :\ s\equiv i\pmod{n}}(v_{s}^{\delta}+h_{s}^{\delta}).$
We chose $k$ large enough such that $\frac{n(k+1)}{2}>g+\delta$, so by Lemma 2.6, the above mapping cone is quasi-isomorphic to the mapping cone of
$\sum_{s\ :\ s\equiv i\pmod{n}}(v_{s}^{\delta}+h_{s}^{\delta}):\bigoplus_{s\ :\ s\equiv i\pmod{n}}A_{s}^{\delta}\to\bigoplus_{s\ :\ s\equiv i\pmod{n}}B^{\delta},$
which by definition is ${\mathbb{X}}_{i,\mathfrak{t}}^{\delta}(n)$.
Therefore, we have the quasi-isomorphism
$CF^{\delta}(Y_{n}(K),i,\mathfrak{t})\cong{\mathbb{X}}_{i,\mathfrak{t}}^{\delta}(n)$
for all $\delta\geq 0$. Applying [OS08, Lemma 2.7], we conclude that
$CF^{+}(Y_{n}(K),i,\mathfrak{t})\cong{\mathbb{X}}_{i,\mathfrak{t}}^{+}(n).$
Proof of Theorem 3.3 for negative surgery coefficients.
Continue to assume $n>0$, only we consider $Y_{-n}(K)$. Again, we first work with the $\delta$ version for $\delta\geq 0$. In this case, let $W^{\prime}_{n(k-1)}(K)$ be the cobordism from $Y_{n(k-1)}(K)$ to $Y$. For each $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$, $i\in{\mathbb{Z}}/n{\mathbb{Z}}$,
$HF^{\delta}(Y_{-n}(K),i,\mathfrak{t})\cong H_{*}\left(MC\left(f_{2,i,\mathfrak{t}}^{\delta}:CF^{\delta}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to\bigoplus^{nk}CF^{\delta}(Y,\mathfrak{t})\right)\right),$
where
$f_{2,i,\mathfrak{t}}^{\delta}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k-1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}:CF^{\delta}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to CF^{\delta}(Y,\mathfrak{t}),$
$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k-1)}(K))\$$|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k-1)\equiv 2s\pmod{2n(k-1)}$,
$\text{and }s\equiv i\pmod{n}\}$,
$CF^{\delta}(Y_{n(k-1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k-1){\mathbb{Z}}\ |\ s\equiv i\pmod{n}}CF^{\delta}(Y_{n(k-1)}(K),s,\mathfrak{t}).$
Consider the summand $CF^{\delta}(Y_{n(k-1)}(K),s,\mathfrak{t})$ for
$\frac{n(k-1)}{2}\leq s<\frac{n(k-1)}{2},\hskip 14.22636pts\equiv i\pmod{n}$.
Then, as above, if $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then $F_{W^{\prime}_{n(k-1)}(K),\mathfrak{s}_{s}}^{\delta}=0$. The conclusion follows again by the same argument as in the positive surgery case.
4 Rational Surgeries
In this section, we sketch the proof of the rational surgery formula:
Theorem 4.1.
Let $K\subset Y$ be a null-homologous knot, and let $p,q$ be a pair of relatively prime integers. Then, for each $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$, $i\in{\mathbb{Z}}/p{\mathbb{Z}}$, there is an isomorphism of ${\mathbb{Z}}[U]$-modules
$H_{*}(\mathbb{X}_{p/q,i,\mathfrak{t}}^{+})\cong HF^{+}(Y_{p/q}(K),i,\mathfrak{t}).$
Ozsváth and Szabó stated and proved this theorem in the case where $\mathfrak{t}$ has torsion first Chern class [OS10]. In this case, they showed this isomorphism also preserves $\mathbb{Z}$}-grading. Their argument still holds when the $\operatorname{Spin}^{c}$ structure is nontorsion, as long as grading information is ignored.
Let $a=\lfloor p/q\rfloor$ and $p/q=a+r/q$. Consider the Hopf link in $S^{3}$ and perform q/r-q/r surgery on one component of the link. Let $O_{q/r}\subset-L(q,r)$ be the other component of the link viewed as a knot in the resulting lens space after surgery. By Kirby calculus, $Y_{p/q}(K)$ can be realized by surgery with coefficient $a$ inside the knot $K\#O_{q/r}\subset Y\#(-L(q,r))$ [OS10]. Thus, one obtains the rational surgery formula by proving a formula for Morse surgeries on rationally null-homologous knots and then applying a Kunneth principle for connected sums.
Analogously to the case of integer surgeries on null-homologous knots, for $K\subset Y$ a rationally null-homologous knot with an orientation (we use $\overline{K}$ to denote the oriented knot), $\lambda$ a framing on $K$, and $\mathfrak{s}\in\operatorname{Spin}^{c}(Y_{\lambda}(K))$, define the chain map $D_{\mathfrak{s}}^{+}:{\mathbb{A}}_{\mathfrak{s}}^{+}(Y,\overline{K})\to{\mathbb{B}}_{\mathfrak{s}}^{+}(Y,\overline{K})$ as in [OS10] Theorem 6.1. Let ${\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda)$ be the mapping cone of $D_{\mathfrak{s}}^{+}$ [OS10].
Theorem 4.2 (Morse surgery formula [OS10] Theorem 6.1).
Let $K\subset Y$ be a rationally null-homologous knot, and let $\lambda$ be a framing on $K$. Let $\mathfrak{s}\in\operatorname{Spin}^{c}(Y_{\lambda}(K))$, then
$HF^{+}(Y_{\lambda}(K),\mathfrak{s})\cong H_{*}({\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda))$
as ${\mathbb{Z}}[U]$-modules.
We recall the definition of $U$-knots from [OS10] Definition 5.2.
Definition 4.3.
[…]
Proposition 4.4 ([OS10], Corollary 5.3).
If $K_{2}\subset Y_{2}$ is a $U$-knot, then for each $\xi_{1}\in\underline{\operatorname{Spin}}^{c}(Y_{1},\overline{K}_{1})$ and $\mathfrak{s}_{2}\in\operatorname{Spin}^{c}(Y_{2})$, there is some $\xi_{2}\in\underline{\operatorname{Spin}^{c}}(Y_{2})$ representing $\mathfrak{s}_{2}$, with the property that
$CFK^{\infty}(Y_{1},\overline{K_{1}},\xi_{1})\cong CFK^{\infty}(Y_{1}\#Y_{2},\overline{K_{1}}\#\overline{K_{2}},\xi_{1}\#\xi_{2})$
[…]
Lemma 4.5 ([OS10], Lemma 7.1).
$O_{q/r}\subset-L(q,r)$ is a $U$-knot. There is an affine identification $\phi$ of the $\operatorname{Spin}^{c}$ structures on $\widehat{HFK}(L(q,r),O_{q/r})$ with $\mathbb{Z}$} such that
$\widehat{HFK}(L(q,r),O_{q/r},\phi(i))\cong\begin{cases}{\mathbb{Z}}&\text{if $0\leq i\leq q-1$,}\\ 0&\text{otherwise.}\end{cases}$
Let $K\subset Y$ be a null-homologous knot. As in section 2.2 of [OS10], $H^{2}(Y,K)$ is defined to be $H^{2}(Y-\operatorname{nd}(K),\partial\operatorname{nd}(K))$ and is an affine space for $\underline{\operatorname{Spin}^{c}}(Y,K)$. There is a natural map $\underline{\operatorname{Spin}^{c}}(Y,K)\to\operatorname{Spin}^{c}(Y)$ which allows us to identify
$\operatorname{Spin}^{c}(Y)\cong\frac{\underline{\operatorname{Spin}^{c}}(Y,K)}{{\mathbb{Z}}\cdot\operatorname{PD}[\mu]}.$
Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, and under the above identification, let $H^{2}(Y,K,\mathfrak{t})\cong{\mathbb{Z}}$ denote the subgroup of $H^{2}(Y,K)$ corresponding to $\mathfrak{t}$, and also $H^{2}(Y\#-L(q,r),K\#O_{q,r},\mathfrak{t})\cong{\mathbb{Z}}$.
With this notation, the following more general version of [OS10, Lemma 7.2] holds with analogous proof.
Lemma 4.6 ([OS10], Lemma 7.2).
Let $K\subset Y$ be a null-homologous knot. Under the connected sum
$(Y,K)\#(-L(q,r),O_{q,r})\to(Y\#-L(q,r),K\#O_{q,r}),$
the diagram in Figure 2 commutes where $f(x,y)=qx+y$. Moreover, under this correspondence, if $K_{\lambda}$ is the pushoff of $K$ with respect to the integral framing $a$, then $\operatorname{PD}[K_{\lambda}]$ represents the element $p\in{\mathbb{Z}}\cong H^{2}(Y\#-L(q,r),K\#O_{q,r})$, where $p/q=a+r/q$.
$\mathbb{Z}\oplus\mathbb{Z}$$\mathbb{Z}$$H^{2}(Y,K,\mathfrak{t})\oplus H^{2}(-L(q,r),O_{q,r})$$H^{2}(Y\#-L(q,r),K\#O_{q,r},\mathfrak{t})$$\scriptstyle{f}\scriptstyle{\cong}\scriptstyle{\cong}$
Figure 2:
Proof.
Follows as in [OS10]. ∎
With these ingredients, the proof of Theorem 4.1 follows as in [OS10].
5 Cosmetic Surgeries
In this section, we prove our main results. For simplicity, we consider coefficients in the field ${\mathbb{F}}={\mathbb{F}}_{2}$. Throughout, we let $K\subset Y$ be a null-homologous knot in a closed, oriented three-manifold. We will obstruct the result of Dehn surgeries on $K$ from being homeomorphic to each other using the rank of $\widehat{HF}(Y_{r}(K))$ for various positive rational numbers $r$.
We state some lemmas that will be useful in calculating the homology of the mapping cones for Dehn surgery.
The following lemma was proven in [NI20] Equations (5) and (6).
Lemma 5.1.
If $s\leq t$, then $\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{t})_{*}$ and $\operatorname{Im}(\widehat{h}_{s})_{*}\supset\operatorname{Im}(\widehat{h}_{t})_{*}$.
As a consequence of the hat version of Lemma 2.5, $\operatorname{rank}H_{*}(\widehat{A}_{s})=\operatorname{rank}H_{*}(\widehat{A}_{-s})$ and $\operatorname{rank}(\widehat{v}_{s})_{*}=\operatorname{rank}(\widehat{h}_{-s})_{*}$ for all $s$.
Lemma 5.2.
Suppose for some $s\in{\mathbb{Z}}$ that $\widehat{v}_{s}$ or $\widehat{h}_{s}$ induces an isomorphism on homology. Then $v_{s}^{+}$ or $h_{s}^{+}$ also induces an isomorphism on homology, respectively.
Proof.
We prove this lemma for $\widehat{v}_{s}$, and the case for $\widehat{h}_{s}$ is proven similarly. We will show by induction that $(v_{s}^{\delta})_{*}$ is an isomorphism for all $\delta\geq 0$, which by Lemma 2.7 in [OS08] implies that $(v_{g-1}^{+})_{*}$ is an isomorphism. Suppose that $(v_{s}^{\delta})_{*}$ is an isomorphism for some $\delta\geq 0$ (the base case $\delta=0$ is $v_{s}^{0}=\widehat{v}_{s}$).
Clearly, the maps $\widehat{v}_{s}$, $v_{s}^{\delta+1}$, and $v_{s}^{\delta}$ fit into the commutative diagram of short exact sequences in Figure 3. Therefore, the diagram in Figure 4 commutes and the rows are exact, so by the five lemma, $(v_{s}^{\delta+1})_{*}$ is also an isomorphism. ∎
[…]
Figure 3:
$H_{*}(A_{s}^{\delta})$$H_{*}(\widehat{A}_{s})$$H_{*}(A_{s}^{\delta+1})$$H_{*}(A_{s}^{\delta})$$H_{*}(\widehat{A}_{s})$$H_{*}(B^{\delta})$${H_{*}(\widehat{B})}$$H_{*}(B^{\delta+1})$$H_{*}(B^{\delta})$${H_{*}(\widehat{B})}\scriptstyle{\partial}$$\scriptstyle{(v_{s}^{\delta})_{*}}$$\scriptstyle{i_{*}}$$\scriptstyle{(\widehat{v}_{s})_{*}}$$\scriptstyle{U_{*}}$$\scriptstyle{(v_{s}^{\delta+1})_{*}}\scriptstyle{\partial}$$\scriptstyle{(v_{s}^{\delta})_{*}}$$\scriptstyle{(\widehat{v}_{s})_{*}}\scriptstyle{\partial}$$\scriptstyle{i_{*}}$$\scriptstyle{U_{*}}\scriptstyle{\partial}$
Figure 4:
Lemma 5.3 ([NI20] Lemma 4.1).
Let $K\subset Y$ be a null-homologous knot. Suppose that $F$ is a taut Seifert surface for $K$. Let $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ for some $N$, and let $\iota:(Y,K)\to(X,K\times\{0\})$ be the inclusion map. Let $G\subset X$ be a properly-embedded smooth connected surface with $\partial G=K\times\{0\}$, $G]=\iota_{*}[F]\in H^{2}(X,K\times\{0\})$, and $g(G)<g(F)$. Then
$\operatorname{Im}(v_{k}^{+})_{*}\supset\operatorname{Im}(h_{k}^{+})_{*},$
when $k\geq g(G)$. Moreover, if $G$ is a disk, then
$(v_{0}^{+})_{*}=(h_{0}^{+})_{*}$
The same is true for the hat version of knot Floer homology by essentially the same proof.
Again, call a knot where such $(X,G)$ for $G$ a disk exists a blow-up bounding knot. We also consider the unknot itself to be a blow-up bounding knot.
The following argument is due to Ni:
Lemma 5.4.
[…]
Proof.
Denote by $O$ the unknot in $Y$. There is a sequence of $N$ crossing changes changing a positive crossing to a negative one that changes $O$ to the knot $K$. Let $W_{n}^{\prime}(K):Y_{n}(K)\to Y$ and $W_{n}^{\prime}(O):Y_{n}(O)\to Y$ denote the natural two-handle cobordisms. Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ on $Y$, let $\mathfrak{x}_{0}(O)\in\operatorname{Spin}^{c}(W_{n}^{\prime}(O))$ and $\mathfrak{x}_{0}(K)\in\operatorname{Spin}^{c}(W_{n}^{\prime}(K))$ as in Section 2 restrict to the $\operatorname{Spin}^{c}$ strctures $\mathfrak{x}_{0}(O)|Y=\mathfrak{x}_{0}(K)|Y=\mathfrak{t}$ and $\mathfrak{t}_{0}(O)=\mathfrak{x}_{0}(O)|Y_{n}(O)$, $\mathfrak{t}_{0}(K)=\mathfrak{x}_{0}(K)|Y_{n}(K)$. Let $\mathfrak{S}$ be the $\operatorname{Spin}^{c}$ structure on $W_{n}^{\prime}(O)\#N\overline{{\mathbb{C}}P^{2}}$ given by the connected sum of $\mathfrak{x}_{0}(O)$ with $N$ copies of $\mathfrak{s}_{1}\in\operatorname{Spin}^{c}(\overline{{\mathbb{C}}P^{2}})$, where $\mathfrak{s}_{1}$ satisfies that $c_{1}(\mathfrak{s}_{1})$ generates $H^{2}(\overline{{\mathbb{C}}P^{2}})$.
By Kirby calculus, the Kirby diagram of the $n$-framed unknot in $Y$ blown up $N$ times is equivalent to the diagram of the $n$-framed knot $K$ in $Y$. Thus, the cobordism $W_{n}^{\prime}(O)\#N\overline{{\mathbb{C}}P^{2}}$ is the composition of two cobordisms
$Y_{n}(O)\xrightarrow{X_{n}(K)}Y_{n}(K)\xrightarrow{W_{n}^{\prime}(K)}Y$
for some cobordism $X_{n}(K)$. The $\operatorname{Spin}^{c}$ structure $\mathfrak{S}$ restricts to $\mathfrak{t}_{0}(O)$ on $Y_{n}(O)$, $\mathfrak{t}_{0}(K)$ on $Y_{n}(K)$, and $\mathfrak{t}$ on $Y$. Therefore, $X_{n}(K)$ induces a cobordism map
$\Phi=\widehat{F}_{X_{n}(K),\mathfrak{S}|X_{n}(K)}:\widehat{HF}(Y_{n}(O),\mathfrak{t}_{0}(O))\to\widehat{HF}(Y_{n}(K),\mathfrak{t}_{0}(K))$
that by Theorem 2.7 fits into the commutative diagram in Figure 5. By a sample calculation, $(\widehat{v}_{0}(O))_{*}$ is surjective. Therefore $(\widehat{v}_{0}(K))_{*}$ must also be surjective. Direct summing over all $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ proves the statement.
$\widehat{HF}(Y_{n}(O),\mathfrak{t}_{0}(O))$$\widehat{HF}(Y_{n}(K),\mathfrak{t}_{0}(K))$$H_{*}(\widehat{A}_{0,\mathfrak{t}}(O))$$H_{*}(\widehat{A}_{0,\mathfrak{t}}(K))$$H_{*}(\widehat{B}_{\mathfrak{t}}(O))$$H_{*}(\widehat{B}_{\mathfrak{t}}(K))$$\scriptstyle{\Phi}\scriptstyle{\cong}\scriptstyle{\cong}$$\scriptstyle{(\widehat{v}_{0}(O))_{*}}$$\scriptstyle{(\widehat{v}_{0}(K))_{*}}$$\scriptstyle{=}$
Figure 5:
The following result and its proof are analogous to [OS10] Proposition 9.7.
Proposition 5.5.
Let $K\subset Y$ be a null-homologous knot. Then
$g(K)=\max(0,\{s\in{\mathbb{Z}}\ |\ (\widehat{v}_{s-1})_{*}\text{ is not an isomorphism}\}).$
In particular, $K$ is the unknot if and only if $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$.
Proof.
Let $g=g(K)$, by Lemma 2.6,
$g(K)\geq\max(0,\{s\in{\mathbb{Z}}\ |\ (\widehat{v}_{s-1})_{*}\text{ is not an isomorphism}\}).$
If $g(K)=0$, i.e. $K$ is the unknot, then $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$. Otherwise, if $g(K)>0$, it suffices to show $(\widehat{v}_{g-1})_{*}$ is not an isomorphism. Consider the short exact sequence
$0\to C\{i<0,j\geq g-1\}\to C\{i\geq 0\text{ or }j\geq g-1\}\xrightarrow{v_{g-1}^{+}}C\{i\geq 0\}\to 0.$ (5.6)
The middle term is $A_{g-1}^{+}$, and $C\{i\geq 0\}=B^{+}$. The assumption that $C$ is reduced together with the $U$-isomorphism and the fact that knot Floer homology detects genus implies that $C\{i,j\}$ is nonzero only if $g\leq i-j\leq g$. Therefore,
$H_{*}(C\{i<0,j\geq g-1\})\cong H_{*}(C\{(-1,g-1)\}),$
which by the $U$-isomorphism is isomorphic to
$H_{*}(C\{0,g\})=\widehat{HFK}(K,g),$
and $\widehat{HFK}(K,g)\neq 0$ since knot Floer homology detects genus. Therefore, $(v_{g-1}^{+})_{*}$ is not an isomorphism. By Lemma 5.2, $(\widehat{v}_{g-1})_{*}$ is also not an isomorphism.
We abuse notation to let $\widehat{A}_{j/q}$ denote $\widehat{A}_{\lfloor j/q\rfloor}$, and also $\widehat{B}_{j/q}$, $\widehat{v}_{j/q}$, $\widehat{h}_{j/q}$ similarly.
Lemma 5.7.
[…]
Proof.
Since $K$ is blow-up bounding, by Lemma 5.3, for $s\geq 0$, we have $\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s+1})_{*}\subset\dots$, and for $s\leq 0$, $\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s-1})_{*}\subset\dots$. Thus, in the first case, there exists $y_{1}\in H_{*}(\widehat{A}_{(j+p)/q})$ such that $(\widehat{h}_{j/q})_{*}(x)=(\widehat{v}_{(j+p)/q})_{*}(y_{1})$. If $(\widehat{h}_{(j+p)/q})_{*}(y_{1})=0$, then stop and set $\alpha=y_{1}$, otherwise keep going. Inductively, given $y_{n}$ in $H_{*}(\widehat{A}_{(j+np)/q})$, there exists $y_{n+1}\in H_{*}(\widehat{A}_{(j+(n+1)p)/q})$ such that $(\widehat{h}_{(j+np)/q})_{*}(y_{n})=(\widehat{v}_{(j+(n+1)p)/q})_{*}$. By Lemma 2.6, this process eventually terminates and gives an element $\alpha=\sum_{n\geq 1}y_{n}\in H_{*}(\oplus_{i\geq j+p}\widehat{A}_{i/q})$ with the desired condition.
Similarly, in the case where $\lfloor j/q\rfloor\leq 0$, since $\operatorname{Im}(\widehat{v}_{j/q})_{*}\subset\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, we can inductively add cancellation terms to the left. ∎
Using the mapping cone formula and the above lemmas, in the case where $K$ is blow-up bounding, we may now calculate the rank of $\widehat{HF}(Y_{r}(K))$ for any positive rational number $r$.
Lemma 5.8.
Let $K\subset Y$ be a blow-up bounding knot, let $p,q>0$ coprime. Let $b=\operatorname{rank}(H_{*}(\widehat{B}))=\operatorname{rank}\widehat{HF}(Y)$. Then
$\begin{split}\operatorname{rank}\widehat{HF}(Y_{p/q}(K))&=q(\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*})\\ &+2q\sum_{s=1}^{g-1}(\operatorname{rank}\ker(\widehat{v}_{s})_{*}+b-\operatorname{rank}(\widehat{v}_{s})_{*})+2t_{K}^{p/q}-pb,\end{split}$ (5.9)
where
$t_{K}^{p/q}$ $=\sum_{j=0}^{p-1}\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right).$ (5.10)
Note that if $Y=S^{3}$, Equation 5.9 agrees with Equation 40 in [OS10]. Indeed, define $\nu(K)$ as in [OS10]. If $\nu(K)>0$, then
$\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right)=\begin{cases}1&j\geq q\nu(K)\text{ and }j<p-q(\nu(K)-1)\\ 0&\text{otherwise.}\end{cases}$
Thus,
$t_{K}^{p/q}=\max(0,p-(2\nu(K)-1)q).$
Also, for all $s\geq 0$,
$\operatorname{rank}\ker(\widehat{v}_{s})_{*}=\begin{cases}\operatorname{rank}H_{*}(\widehat{A}_{s})_{*}&s<\nu(K)\\ \operatorname{rank}H_{*}(\widehat{A}_{s})_{*}-1&s\geq\nu(K),\end{cases}$
so the two equations agree.
If $\nu(K)=0$, then $t_{K}^{p/q}=p$ and $\operatorname{rank}\ker(\widehat{v}_{s})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{s})_{*}-1$ for all $s\geq 0$, so again the two equations agree.
Proof.
By Theorem 4.1, $\widehat{HF}(Y_{p/q}(K))\cong H_{*}(\widehat{\mathbb{X}}_{p/q})$, the homology of the mapping cone of $\widehat{D}_{p/q}$ from $\widehat{{\mathbb{A}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{A}}_{i}$ to $\widehat{{\mathbb{B}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{B}}_{i}$ (we have omitted the slope $p/q$ from the notation and summed over all $\operatorname{Spin}^{c}$ structures on $Y$). So
$\widehat{\mathbb{A}}=\bigoplus_{i=0}^{p-1}\prod_{r\in{\mathbb{Z}}}\widehat{A}_{(i+pr)/q}=\prod_{j\in{\mathbb{Z}}}\widehat{A}_{j/q}$
has $q$ copies of $\widehat{A}_{s}$ for each $s\in{\mathbb{Z}}$, and under the map $\widehat{v}$, each $\widehat{A}_{j/q}$ corresponds to a copy of $\widehat{B}$ in $\widehat{\mathbb{B}}$, which we shall denote $\widehat{B}_{j/q}$. Then $\widehat{h}_{j/q}$ maps $\widehat{A}_{j/q}$ to $\widehat{B}_{(j+p)/q}$. Thus, for each $s\in{\mathbb{Z}}$, there are $p$ copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s-1}$ and qpq-p copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s}$.
The finiteness of $H_{*}(\widehat{\mathbb{X}}_{p/q})$ is guaranteed by Lemma 2.6; in fact, it implies that $\widehat{\mathbb{X}}_{p/q}$ is isomorphic to a truncated version of the mapping cone
$\widehat{\mathbb{X}}_{p/q}^{c}=MC(\widehat{D}_{p/q}^{c}:\widehat{\mathbb{A}}_{p/q}^{c}\to\widehat{\mathbb{B}}_{p/q}^{c}),$
for some large $c$, where
$\widehat{\mathbb{A}}_{p/q}^{c}=\bigoplus_{j=-qc+1}^{qc-1}\widehat{A}_{j/q}=\bigoplus_{s=-c+1}^{c-1}\bigoplus_{i=0}^{q-1}\widehat{A}_{s},\hskip 14.22636pt\widehat{\mathbb{B}}_{p/q}^{c}=\bigoplus_{j=-qc+p+1}^{qc-1}\widehat{B}_{j/q},$
and $\widehat{D}_{p/q}^{c}$ is the restriction of $\widehat{D}_{p/q}$ to $\widehat{A}_{p/q}^{c}$, with the exception that for any $x\in\widehat{A}_{j/q}$ with $qc+1\leq j\leq-qc+p$, $\widehat{D}_{p/q}^{c}(x)=\widehat{h}_{j/q}(x)$, and for any $x\in\widehat{A}_{j/q}$ with $qc-p\leq j\leq qc-1$, $\widehat{D}_{p/q}^{c}(x)=\widehat{v}_{j/q}(x)$. Note that it suffices to take $c\geq g+p/q+1$, and
$\operatorname{rank}H_{*}(\widehat{\mathbb{A}}_{p/q}^{c})=q\sum_{s=-c+1}^{c-1}\operatorname{rank}H_{*}(\widehat{A}_{s}),\hskip 14.22636pt\operatorname{rank}H_{*}(\widehat{\mathbb{B}}_{p/q}^{c})=qb(2c-1)-pb.$
We construct a basis for $\ker(\widehat{D}_{p/q})_{*}$ of the entire mapping cone, which has the same rank as $\ker(\widehat{D}_{p/q}^{c})_{*}$ for large $c$ by Lemma 2.6. For $j\geq 0$, let
$\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$
be a basis of $\ker(\widehat{v}_{j/q})_{*}$. For each such $x_{j/q}^{i}$, choose an $\alpha_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\alpha_{j/q}^{i}$ is in $\ker(\widehat{D}_{p/q})_{*}$. For $j<0$, let
$\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{h}_{j/q})_{*}}$
be a basis of $\ker(\widehat{h}_{j/q})_{*}$. For each such $x_{j/q}^{i}$, choose a $\beta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\beta_{j/q}^{i}$ is in $\ker(\widehat{D}_{p/q})_{*}$. Lastly, for $0\leq j\leq p-1$, let
$\{y_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right)}$
be a basis of a subspace of $H_{*}(\widehat{A}_{j/q})$ that is mapped isomorphically to $\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$ under $(\widehat{v}_{j/q})_{*}$. Choose $z_{j/q}^{i}\in H_{*}(\widehat{A}_{(j-p)/q})$ such that $(\widehat{v}_{j/q})_{*}(y_{j/q}^{i})=(\widehat{h}_{(j-p)/q})_{*}(z_{j/q}^{i})$. Choose $\gamma_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ and $\delta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{y}_{j/q}^{i}=y_{j/q}^{i}+z_{j/q}^{i}+\gamma_{j/q}^{i}+\delta_{j/q}^{i}$ is in $\ker(\widehat{D}_{p/q})_{*}$.
The set
$S=\{\widetilde{x}_{j/q}^{i}\}\cup\{\widetilde{y}_{j/q}^{i}\}$
is linearly independent, and it is finite by Lemma 2.6. Moreover, it spans $\ker(\widehat{D}_{p/q})_{*}$: let $z\in\ker(\widehat{D}_{p/q})_{*}$ be nonzero, and suppose first that $z\in H_{*}(\oplus_{k\geq 0}\widehat{A}_{k/q})$. Let $z_{j/q}$ denote the term of $z$ in $H_{*}(\widehat{A}_{j/q})$, and let
$j=\min\{k\geq 0\ |\ z_{k/q}\neq 0\}.$
Then $z\in\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$, so $z_{j/q}\in\operatorname{Span}\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$: otherwise if $z_{j/q}\notin\ker(\widehat{v}_{j/q})_{*}$, then since $z\in H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$, we have $z_{(j-p)/q}=0$ so
$(\widehat{v}_{j/q})_{*}(z_{j/q})+(\widehat{h}_{(j-p)/q})_{*}(z_{(j-p)/q})\neq 0\in H_{*}(\widehat{B}_{j/q})$
which means $z\notin\ker(\widehat{D}_{p/q})_{*}$. Thus, we can subtract some combination of the basis elements $\widetilde{x}_{j/q}^{i}$ to eliminate the $H_{*}(\widehat{A}_{j/q})$ component of $z$:
$z-\sum_{i\in J}\widetilde{x}_{j/q}^{i}\in\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k>j}\widehat{A}_{k/q})$
for some indexing set $J\subset\{1,\dots,\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}\}$. Inductively, we can then write
$z=\sum_{k\in I}\sum_{i_{k}\in J_{k}}\widetilde{x}_{k/q}^{i_{k}}$
for some finite indexing sets $I\subset\{k\geq j\}$, $J_{k}\subset\{1,\dots,\operatorname{rank}\ker(\widehat{v}_{k/q})_{*}\}$. Similarly, if $z\in H_{*}(\oplus_{k<0}\widehat{A}_{k/q})$, then we can write
$z=\sum_{k\in I}\sum_{i_{k}\in J_{k}}\widetilde{x}_{k/q}^{i_{k}}$
for some finite indexing set $I\subset\{k\leq j\}$, $J_{k}\subset\{1,\dots,\operatorname{rank}\ker(\widehat{h}_{k/q})_{*}\}$.
Now for the general case $z\in\ker(\widehat{D}_{p/q})_{*}$, we claim that for each $0\leq j\leq p-1$,
$z_{j/q}\in\operatorname{Span}\left(\{y_{j/q}^{i}\}\cup\{x_{j/q}^{i}\}\right).$
Indeed, suppose this is not true for some $0\leq j\leq p-1$, then $(\widehat{v}_{j/q})_{*}(z_{j/q})\notin\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, so $(\widehat{D}_{p/q})_{*}(z)\neq 0$, a contradiction. Thus,
$z-\sum_{j=0}^{p-1}\left(\sum_{i\in J_{j}}\widetilde{y}_{j/q}^{i}+\sum_{i\in J_{j}^{\prime}}\widetilde{x}_{j/q}^{i}\right)$ $\in\left(\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k\geq p}\widehat{A}_{k/q})\right)$
$\hskip 14.22636pt+\left(\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k<0}\widehat{A}_{k/q})\right).$
By the above argument, $z$ is in the span of $S$.
Therefore, $S$ is a basis for $\ker(\widehat{D}_{p/q})_{*}$, so since $\operatorname{rank}\ker(\widehat{v}_{s})_{*}=\operatorname{rank}\ker(\widehat{h}_{-s})_{*}$ by the hat version of Lemma 2.5 and these are 0 when $s\geq g$ by Lemma 2.6,
$\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}=q\operatorname{rank}\ker(\widehat{v}_{0})_{*}+2q\sum_{s=1}^{g-1}\operatorname{rank}\ker(\widehat{v}_{s})_{*}+t_{K}^{p/q}.$
Since $\operatorname{rank}(\widehat{v}_{s})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{s})-\operatorname{rank}\ker(\widehat{v}_{s})_{*}$ and $\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}=\operatorname{rank}H_{*}(\widehat{\mathbb{B}}_{p/q}^{c})-\operatorname{rank}H_{*}(\widehat{\mathbb{A}}_{p/q}^{c})+\operatorname{rank}\ker(\widehat{D}_{p/q}^{c})_{*}$ for some large $c$, it follows that
$\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}$ $=b(2c-1)q-pb-q\sum_{s=-c+1}^{c-1}\operatorname{rank}H_{*}(\widehat{A}_{s})$
$\hskip 14.22636pt+q\operatorname{rank}\ker(\widehat{v}_{0})_{*}+2q\sum_{s=1}^{g-1}\operatorname{rank}\ker(\widehat{v}_{s})_{*}+t_{K}^{p/q}$
$=q(b-\operatorname{rank}H_{*}(\widehat{A}_{0})+\operatorname{rank}\ker(\widehat{v}_{0})_{*})$
$\hskip 14.22636pt+2q\sum_{s=1}^{g-1}(b-\operatorname{rank}H_{*}(\widehat{A}_{s})+\operatorname{rank}\ker(\widehat{v}_{s})_{*})+t_{K}^{p/q}-pb$
$=q(b-\operatorname{rank}(\widehat{v}_{0})_{*})+2q\sum_{s=1}^{g-1}(b-\operatorname{rank}(\widehat{v}_{s})_{*})+t_{K}^{p/q}-pb,$
where the second equality uses that $\operatorname{rank}H_{*}(\widehat{A}_{s})=b$ for all $|s|\geq g$ by Lemmas 2.5 and  2.6. Therefore, because
$\operatorname{rank}(H_{*}(\widehat{\mathbb{X}}_{p/q}))=\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}+\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q})_{*}$
and
$\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}=\operatorname{rank}\ker(\widehat{D}_{p/q}^{c})_{*},\hskip 14.22636pt\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q})_{*}=\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}$
for large enough $c$, equation 5.9 holds.
Now we are ready to prove our main theorems. The main ingredient for each result is Equation 5.9.
Theorem 5.11.
Let $K\subset Y$ be a blow-up bounding knot, and suppose $0<r<s\leq 1$ are distinct positive rational numbers with the property that $Y_{r}(K)\cong\pm Y_{s}(K)$. Then $K$ is the unknot. Furthermore, if $K\subset Y$ is a blow-up unknotted knot and $0<r<s$ are distinct positive rational numbers such that $Y_{r}(K)\cong\pm Y_{s}(K)$, then $r,s>1$ or $K$ is the unknot.
Proof.
Let $K\subset Y$ be blow-up bounding. Since $H_{1}(Y_{\pm p/q}(K))\cong H_{1}(Y)\oplus{\mathbb{Z}}/p{\mathbb{Z}},$ we can fix $p>0$ throughout. First, let $q$ be relatively prime to $p$ and $q\geq p$. Then $\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}=\operatorname{Im}(\widehat{v}_{0})_{*}$ for all $j=0,\dots,p-1$. Thus $t_{K}^{p/q}=p\operatorname{rank}(\widehat{v}_{0})_{*}$. Therefore, according to equation 5.9, for positive integral $q$ with $0<p/q\leq 1$, $\operatorname{rank}\widehat{HF}(Y_{p/q}(K))$ is a nondecreasing function of $q$. The function is strictly increasing unless $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$. Hence, by Proposition 5.5, $K$ is the unknot.
Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the first part of the theorem is true. Also, the second part is true for $0<r<s\leq 1$ since blow-up unknotted knots are blow-up bounding.
Now suppose $K$ is blow-up unknotted and $q,q^{\prime}$ are relatively prime to $p$ and $0<q<p<q^{\prime}$. Then, with $)b=\operatorname{rank}(H_{*}(\widehat{B}$)),
$\begin{split}t_{K}^{p/q}-t_{K}^{p/q^{\prime}}&\leq\sum_{j=0}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-p\operatorname{rank}(\widehat{v}_{0})_{*}\\ &=\sum_{j=q}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-(p-q)\operatorname{rank}(\widehat{v}_{0})_{*}\\ &\leq(p-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*})\\ &=0\end{split}$ (5.12)
since $(\widehat{v}_{0})_{*}$ is surjective by Lemma 5.4. By equation 5.9,
$\operatorname{rank}\widehat{HF}(Y_{p/q^{\prime}}(K))$ $-\operatorname{rank}\widehat{HF}(Y_{p/q}(K))=(q^{\prime}-q)(\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*})$
$+(q^{\prime}-q)\left(2\sum_{s=1}^{g-1}(\operatorname{rank}\ker(\widehat{v}_{s})_{*}+b-\operatorname{rank}(\widehat{v}_{s})_{*})\right)$
$+2t_{K}^{p/q^{\prime}}-2t_{K}^{p/q}.$
The entire sum from $s=1$ to g1g-1 on the right hand side is nonnegative, so we ignore this term, and by Lemma 5.4, we have
$\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*}=\operatorname{rank}\ker(\widehat{v}_{0})_{*}.$
Together with inequality 5.12, we get
$\operatorname{rank}\widehat{HF}(Y_{p/q^{\prime}}(K))-\operatorname{rank}\widehat{HF}(Y_{p/q}(K))$ $\geq(q^{\prime}-q)\operatorname{rank}\ker(\widehat{v}_{0})_{*}$
$\geq 0,$
with equality only if $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$, hence $K$ is the unknot by Lemma 5.5.
Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the result holds. ∎
In the case where $p$ is $1$ or $2$, the following corollary is immediate, since there do not exist distinct positive integers $q,q^{\prime}$ coprime to $p$ with $p/q$ and $p/q^{\prime}$ both greater than $1$.
Corollary 5.13.
Let $K\subset Y$ be a blow-up unknotted knot, and suppose $0<q^{\prime}<q$ are distinct positive numbers with the property that $Y_{p/q}(K)\cong\pm Y_{p/q^{\prime}}(K)$ for $p\in\{1,2\}$, then $K$ is the unknot.
Theorem 5.14.
Let $K\subset Y$ be a nontrivial blow-up bounding knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $K$ is a nontrivial blow-up unknotted knot, then $Y_{1}(K)\ncong\pm Y$.
Proof.
Let $q\geq 2$, then by Equation 5.9,
$\operatorname{rank}\widehat{HF}(Y_{1/q}(K))\geq(b-\operatorname{rank}(\widehat{v}_{0})_{*})(q-1)+\operatorname{rank}(\widehat{v}_{0})_{*}+q\operatorname{rank}\ker(\widehat{v}_{0})_{*}\geq b,$
and it is strictly greater unless $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$. Hence, by Proposition 5.5, $K$ is the unknot.
If $q=1$, then
$\operatorname{rank}\widehat{HF}(Y_{1}(K))\geq\operatorname{rank}(\widehat{v}_{0})_{*}+\operatorname{rank}\ker(\widehat{v}_{0})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{0})\geq b$
by Lemma 5.4 since $K$ is blow-up unknotted, and it is strictly greater unless $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$, which implies $K$ is the unknot.
Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the result holds. ∎
Acknowledgements The author would like to thank Yi Ni for advising him in this project and the referee for many helpful comments.